After sampling 16 members of a normal population you find the sample mean equals 50 and the sample standard deviation, s, equals 10. Construct a 95% confidence interval estimate for mu, the population mean.

A. 45.62 – 54.38

B. 44.67 – 55.33

C. 45.10 – 54.90

D. None of the above
Answer #1

Solution :

Given that,

\bar x = 50

s =10

n = Degrees of freedom = df = n - 1 =16 - 1 = 15

At 95% confidence level the t is ,

\alpha = 1 - 95% = 1 - 0.95 = 0.05

  \alpha/ 2= 0.05 / 2 = 0.025

t\alpha /2,df = t0.025,15 = 2.131 ( using student t table)

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 2.131 * ( 10/ \sqrt 16)

= 5.33

The 95% confidence interval mean is,

\bar x - E < \mu < \bar x + E

50 - 5.33 < \mu <50 + 5.33

44.67 – 55.33

(44.67 – 55.33 )